# Amuro & Char: Analyzing the Relationship between Pre-Training and Fine-Tuning of Large Language Models

### v2

@@ Acknowledgments
The authors thank Saleh Soltan, Niyati Bafna, Fan Bai, Miriam Wanner, Xinbo Wu, Carlos Aguirre for their helpful feedback.
Appendix A Hyperparameter Tuning
For both supervised fine-tuning and instruction tuning, we pre-set the effective batch size to 8, and tune the learning rate within $2\times 10^{-5}$, $2\times 10^{-6}$, $2\times 10^{-7}$}. Each model is fine-tuned for 3 epochs on the supervised fine-tuning tasks and 5 epochs on Tulu for instruction tuning. In both settings, we adopt an AdamW optimizer with a linear learning rate scheduler. The optimizer is warmed up for the first $3\%$ of the training time.
Appendix B Prediction Generation Method
@@
For classification tasks, we examine three different prediction generation methods: Free Generation (Free), Constrained Generation (Constrained), and Token Probability (TokenProb), the results are shown in Table 2. In Constrained, we force the output to include at least one label in the acceptable label set. In TokenProb, we compare the logits of acceptable labels, and select the label with the highest score as final output. This ablation study is conducted only on the BASE and fine-tuned versions of final checkpoint of the pre-trained model. We find that, although prediction generation methods bring less effect to the evaluation result of a fine-tuned model, BASE variants suffer much more from not knowing the desired output. Therefore, we proceed all the classification experiments with TokenProb.
B.1 Label and Tokenizations
Depending on the tokenizer variant, the label text may be tokenized differently, leading to evaluation unreliability. For example, in paraphrase detection, the model could assign probability on both “yes" and “ yes" (the same label with a prefix space). This behavior is reported and explored in various related work Sun et al. (2023); Batsuren et al. (2024); Singh and Strouse (2024). In this study, we leniently regard all individual tokens that contains the whole label or part of the label along with some special charcters that do not affect the semantic as an acceptable target label.
Appendix C Task Format
We adopt the task format from Yang et al. (2024), with an additional task format of input-output. How each dataset is formated can be found in Table 4.
Appendix D GPU Hours per-Experiment
We show a table of GPU hours spent for each experiment in Table 3. The total number of GPU hours spent in this project is approximately 1067 A100 hours. We lose track of the GPU hours spent on preliminary experiments, so a lower-bound estimation is reported.
@@ Appendix D GPU Hours per-Experiment
Appendix E Per-dataset Figures
@@
We show the model performance on each dataset after supervised fine-tuning and instruction tuning correspondingly in Figure 9 and Figure 8. The datasets that already show improvement during pre-training do not benefit from fine-tuning, while performance improve drastically on the datasets that the model has never learned during pre-training.
Out-of-domain Generalization
The out-of-domain performance for each dataset with respect to pre-training steps is shown in Figure 10. Overall, the model generalizes well after fine-tuning on NLI tasks, while its performance deteriorates when evaluated on out-of-domain paraphrase detection tasks.
@@
Cross-task Generalization
The cross-task performance for each dataset with respect to pre-training steps is shown in Figure 11 and Figure 12.
@@
Task-Format
The performance of models on evaluation sets formatted with different prompt formatting method is shown in Figure 13.
@@
Appendix F License of Artifacts
We include the license of artifacts used in this paper in Table 5
Name License Name License
OLMo-1b Apache 2.0 SocialIQa CC-BY
TULU ODC-BY CNN/DailyMail Apache 2.0
ARC CC BY-SA TweetQA CC BY-SA-4.0
OpenbookQA Apache 2.0 MNLI CC-BY-3.0
Hellaswag MIT GPT3NLI MIT
BoolQ Apache 2.0 RTE N/A
SciQ CC-BY-NC-3.0 Paws Free
XSum MIT QQP Non-Commercial
XLSum CC-BY-NC-SA 4.0 STS-B Other
Table 5: License of artifacts used in this paper.
Appendix G Performance Difference Numbers
@@
Appendix H Full Performance Table
Due to availability of space and the amount of fine-tuned checkpoints, we omit displaying all exact metric values in the paper. The performance of each fine-tuned variant on each dataset can be found in the csv file in the code base.
Appendix I Generalization Taxonomy
Following the generalization taxonomy in Hupkes et al., 2023, the evaluation card is included in Table I.

### v5

@@ Acknowledgments
The authors thank Saleh Soltan, Niyati Bafna, Fan Bai, Miriam Wanner, Xinbo Wu, and Carlos Aguirre for their helpful feedback. This work was supported in part by a grant from the JHU + Amazon Initiative for Interactive AI.
Appendix A Model and Data Selection
Pythia OpenLLAMA K2 (LLM360) Crystal (LLM360) Baichuan2
Citation Biderman et al.,2023 Geng and Liu,2023 LLM360,2024 Tao et al.,2024 Yang et al.,2023
Size (Param)
70M, 160M,
410M, 1B,
1.4B, 2.8B,
6.9B, 12B
3B, 7B 65B 7B 7B, 13B
Languages English English English English English & Chinese
Pre-trained
Tokens
300B 1T 1.4T 1300B 2.6T
Note - -
Multi-phase
pre-training
Multi-phase
pre-training
OLMO-2 OLMO TinyLLaMA RedPajama-INCITE Bloom
Citation Ai2,2024 Groeneveld et al.,2024 Zhang et al.,2024 Computer,2023 Le Scao et al.,2023
Size (Param) 4T, 5T 1B, 7B 1B 7B 176B
Languages English English English English Multilingual
Pre-trained
Tokens
7B, 13B 3T, 2.5T 3T 1.2T 366B
Note
Multi-phase
pre-training
BOS Token leads
to training history
inconsistency.
Poor
fine-tunablity
Table 2: Large language models with public release of intermediate pre-training checkpoints. All models are under Apache 2.0 license.
Only a small subset of large language models publicly release their intermediate training checkpoints. We list these models in Table 2 and would like to call for model developers to release intermediate checkpoints in the future to aid the research of training dynamics. To reduce the confounding factor of language and stages of training, we select the models that are dominantly trained in English and followed a single-staged training strategy. Only the models pre-trained with significantly more tokens than the model parameters are considered to avoid the occurrence of double descent (Belkin et al., 2019; Schaeffer et al., 2023) in the middle of pre-trianing, which could lead to a broken scaling law (Caballero et al., 2022) that complicates the analysis. Additionally, we restrict our selection to models pre-trained on over one trillion tokens, thereby ensuring a sufficiently extended training trajectory is represented. We conduct initial experiments with OLMo and RedPajama-INCITE. We observe that the RedPajama-INCITE shows subtle improvement following instruction-tuning or fine-tuning, and its 7B variant shows lower performance compared to OLMo 1B. Therefore, we select OLMo 1.0 1B as the backbone for analysis.
During this study, several recent initiatives released the intermediate checkpoints. We also list these works in Table 2.
Appendix B Hyperparameter Tuning
For both supervised fine-tuning and instruction tuning, we pre-set the effective batch size to 8, and tune the learning rate within $2\times 10^{-5}$, $2\times 10^{-6}$, $2\times 10^{-7}$}. OLMo-1B is fine-tuned for 3 epochs on the supervised fine-tuning tasks and 5 epochs on Tulu for instruction tuning. Llama3-8B is fine-tuned for 2 epochs with a learning rate of $5\times 10^{-6}$, with learning rate selected from $5\times 10^{-5}$, $5\times 10^{-6}$, $5\times 10^{-7}$}. In both settings, we adopt an AdamW optimizer with a linear learning rate scheduler. The optimizer is warmed up for the first $3\%$ of the training time.
Appendix C Prediction Generation Method
@@
For classification tasks, we examine three different prediction generation methods: Free Generation (Free), Constrained Generation (Constrained), and Token Probability (TokenProb), the results are shown in Table 3. In Constrained, we force the output to include at least one label in the acceptable label set. In TokenProb, we compare the logits of acceptable labels and select the label with the highest score as the final output. This ablation study is conducted only on the BASE and fine-tuned versions of the final checkpoint of the pre-trained model. We find that, although prediction generation methods have less effect on the evaluation result of a fine-tuned model, BASE variants suffer much more from not knowing the desired output. Therefore, we proceed with all the classification experiments with TokenProb.
C.1 Label and Tokenizations
Depending on the tokenizer variant, the label text may be tokenized differently, leading to evaluation unreliability. For example, in paraphrase detection, the model could assign probability on both “yes" and “ yes" (the same label with a prefix space). This behavior is reported and explored in various related work (Sun et al., 2023; Batsuren et al., 2024; Singh and Strouse, 2024). In this study, we leniently regard all individual tokens that contain the whole label or part of the label along with some special characters that do not affect the semantics as an acceptable target label.
Appendix D Task Format
We adopt the task format from (Yang et al., 2024), with an additional task format of input-output. How each dataset is formated can be found in Table 5.
Appendix E GPU Hours per-Experiment
We show a table of GPU hours spent for each experiment in Table 4. The total number of GPU hours spent on this project is approximately 1067 A100 hours. We lose track of the GPU hours spent on preliminary experiments, so a lower-bound estimation is reported.
@@ Appendix E GPU Hours per-Experiment
Appendix F Per-dataset Figures
@@
We show the model performance on each dataset after supervised fine-tuning and instruction tuning correspondingly in Figure 10 and Figure 9. The datasets that already show improvement during pre-training do not benefit from fine-tuning, while performance improves drastically on the datasets that the model has never learned during pre-training.
Out-of-domain Generalization The out-of-domain performance for each dataset with respect to pre-training steps is shown in Figure 11. Overall, the model generalizes well after fine-tuning on NLI tasks, while its performance deteriorates when evaluated on out-of-domain paraphrase detection tasks.
@@
Cross-task Generalization The cross-task performance for each dataset with respect to pre-training steps is shown in Figure 12 and Figure 13.
@@
Task-Format The performance of models on evaluation sets formatted with different prompt formatting methods is shown in Figure 14.
@@
Appendix G Llama3-8B Results
@@
Figure 15: Ratio of out-of-domain performance change for each task on the final checkpoint of LLAMA3-8B.
To provide more evidence of the findings on a different model architecture and size, we conduct some experiments on the final checkpoint of Llama3-8B.
Task Transfer Similar to our findings with OLMo, Llama3-8B fine-tuned on classification tasks and evaluated on generation tasks decreases on average 61.0% compared to models that are never fine-tuned. In contrast, models fine-tuned on generation tasks perform similarly to the BASE model on classification tasks, with a 10.6% MRC.
Domain Knowledge The ratio of out-of-domain performance change by task is shown in Figure 15. Overall, we observe that Llama and OLMo experience benefits with different tasks after fine-tuning, but both model shows an inconsistent change across tasks.
Name License Name License
OLMo-1b Apache 2.0 SocialIQa CC-BY
TULU ODC-BY CNN/DailyMail Apache 2.0
ARC CC BY-SA TweetQA CC BY-SA-4.0
OpenbookQA Apache 2.0 MNLI CC-BY-3.0
Hellaswag MIT GPT3NLI MIT
BoolQ Apache 2.0 RTE N/A
SciQ CC-BY-NC-3.0 Paws Free
XSum MIT QQP Non-Commercial
XLSum CC-BY-NC-SA 4.0 STS-B Other
Table 7: License of artifacts used in this paper.
Appendix H License of Artifacts
We include the license of artifacts used in this paper in Table 7
Appendix I Full Performance Table
Due to the availability of space and the amount of fine-tuned checkpoints, we omit displaying all exact metric values in the paper. The performance of each fine-tuned variant on each dataset can be found in the csv file under directory results in the code base.
Appendix J Performance Difference Numbers
@@ Appendix I Full Performance Table
Appendix K Generalization Taxonomy
Following the generalization taxonomy in Hupkes et al. (2023), the evaluation card is included in Table 8.
@@
Table 8: Generalization experiment summary following taxonomy in Hupkes et al. (2023).